OpenStudy (anonymous):

Solve the equation by completing the square. 2x^2 + 16x = -14 x= (smaller value) x= (larger value)

OpenStudy (anonymous):

first off if you are going to complete the square best divide by 2. you get \[x^2+8x=-7\]

OpenStudy (anonymous):

then work as follows. half of 8 is 4 so write \[(x+4)^2=-7+4^2\] \[(x+4)^2=9\]

OpenStudy (anonymous):

then take the square root of both sides. dont forget plus and minus \[x+4=3\] or \[x+4=-3\]

OpenStudy (anonymous):

then solve for x to get \[x=-1,x=-7\]

OpenStudy (anonymous):

for your answer it is smaller value is -7, larger is -1